Circular No. 6201

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

EUVE J1429–38.0 N. Craig, University of California at Berkeley, communicates: “The

field of EUVE J1429–38.0 (α = 14^{h}29^{m}24^{s}.0, δ = −38^{o}04′05′′, equinox 2000.0) was observed as part of the EUVE Optical ID Campaign on May 21.169 UT at Cerro Tololo InterAmerican Observatory with the 1.5-m telescope (+ CCD GEC10). The proposed optical counterpart of EUVE J1429– 38.0 (α = 14^{h}29^{m}27^{s}, δ = −38^{o}04′.2, equinox 2000.0) is a star lying 46′′ from the reported EUV center and is well within the source error circle. The estimated apparent magnitude is m_{v} = 12.0. [This may be NSV 06680 = HV 7408—Ed.] Its optical spectrum clearly shows the characteristics of a cataclysmic variable spectrum, with the typical Balmer emission lines, as well as lines of He I, Ca II and strong He II emission. The continuum increases toward the blue at short wavelengths and toward the red at long wavelengths, indicating the presence of a secondary red star. However, the inferred contribution of the companion is not sufficiently strong to allow absorption lines to be detected. EUVE J1429–38.0 was detected serendipitously through the EUVE Right Angle Program (RAP) conducted with the survey (imaging) scanners. As a result of the long exposures typical of spectrometer and Deep Survey observations, the RAP program is up to 20 times more sensitive than the EUVE all-sky survey. The source was detected in the EUVE Lexan/B band on 1993 Mar. 4 with 0.009 cts/s. A further EUVE RAP observation of this previously unidentified source was obtained during 1995 Mar. 22–Apr. 4. These later data sets yield a countrate of 0.001 cts/s, indicating that the source is variable in the EUV. Monitoring photometric and spectroscopic observations are urged. Any detection of optical outbursts should be reported to the EUVE Associate Project Scientist R. Oliversen (oliversen@stars.gsfc.nasa.gov) for possible EUVE TOO (target of opportunity) observations.”

X-RAY NOVA 1994 IN SCORPIUS AND GRS 1915+105 S. Sazonov and R. Sunyaev, on behalf of the GRANAT/WATCH Team,

report: “We are currently observing an outburst in the 1994 Scorpius xray nova GRO J1655–40. The source was first marginally detected by GRANAT/WATCH on July 23 with flux 200 ± 100 mCrab at 8–20 keV. During July 25–30 its intensity daily averages were ∼ 800±200 mCrab. This agrees well with the single power-law spectrum reported by BATSE for the 20–200-keV spectral band (IAUC 6196). The outburst of GRS 1915+105 reported on IAUC 6196 is continuing. On July 29–30 the intensity of the source in the 8–20-keV energy band was as high as 1 ± 0.1 Crab.”

1995 August 4 Brian G. Marsden